Suppose that a machine part requires 400 mg of copper to provide a coating that is thick enough for a certain electrical function. If you run an electrolysis using copper(II) ions for 10 minutes at a current of 0.5 A, will you meet the requirement?

1
Expert's answer
2021-11-24T12:21:56-0500

Current, I = 0.5 A

Time, t = 10 min = 600 s

Quantity of electricity, Q = I × t = (0.5 A) × (600 s) = 300 C

Mass of copper (Cu), m = ??? g

Molar mass of copper (Cu), M = 63.546 g mol-1

Cu2+(aq) + 2e → Cu(s)

Z = 2


Solution:

Calculate the mass of copper produced at an electrode during electrolysis:

Faraday's 1st law can be used.

Faraday's 1st law can be expressed as:

m = (Q × M) / (Z × F)

where:

m = mass of material (g)

M = molar mass of material (g mol-1)

Q = quantity of electricity (C)

Z = number of electrons

F = Faraday constant, 96485 (C mol-1)

 

Thus:

m = (300 C × 63.546 g mol-1) / (2 × 96485 C mol-1) = 0.0988 g Cu = 98.8 mg Cu


98.8 mg < 400 mg

Therefore, we will not meet the requirement.

The produced mass of copper at an electrode is insufficient to provide a coating.


Answer: We will not meet the requirement